OpenStudy (anonymous):

Pilar used four reusable shopping bags on a recent purchase she made at a grocery store. Each bag decreased the amount she spent by 5 cents. What was the change to the amount Pilar spent at the grocery store by using the reusable bags? 20 cents -9 cents -20 cents 9 cents

OpenStudy (mathstudent55):

1 bag reduces 1 * 5 cents = 5 cents 2 bags reduce 2 * 5 cents = 10 cents 3 bags reduce 3 * 5 cents = 15 cents can you continue?